# Robust Multi-Agent Decision-Making in Finite-Population Games

### v1

@@ Abstract
We study the robustness of an agent decision-making model in finite-population games, with a particular focus on the Kullback-Leibler Divergence Regularized Learning (KLD-RL) model. Specifically, we examine how the model’s parameters influence the effects of various sources of noise and modeling inaccuracies – factors commonly encountered in engineering applications of population games – on agents’ decision-making. Our analysis provides insights into how these parameters can be effectively tuned to mitigate such effects. Theoretical results are supported by numerical examples and simulation studies that validate the analysis and illustrate practical strategies for parameter selection.
@@ I Introduction
The population game and evolutionary dynamics framework provides a powerful foundation for modeling and analyzing repeated strategic interactions among a population of decision-making agents [1]. In this work, we consider a class of population games in which agents’ payoffs – driving their strategy selection – are determined by a dynamic payoff mechanism. Specifically, payoff vectors evolve according to a dynamical system model that depends on both the internal state of the game and the population’s state, which reflects agents’ strategy selections. Our primary focus is on a particular subclass of these games, known as task allocation games [2, 3], where agents interact strategically to accomplish a given set of tasks.
The conventional framework relies on several key assumptions: 1)1)1 ) payoff observations are noise-free, 2)2)2 ) switching strategies incurs no delay, and 3)3)3 ) the population size is infinite. However, these assumptions are often unrealistic in engineering applications of population games. We specifically examine decision-making processes in finite-population games, where agents select strategies based on noisy payoff estimates influenced by their own past actions, and where strategy updates are subject to time delays. A central challenge posed by noise and modeling errors is the increased variability in agents’ decision-making, which can hinder convergence to the optimal strategy selection.
@@
This work builds upon and extends recent studies in multi-agent games [5, 6, 7, 8, 9, 10], which examine the effects of noise in payoff observations [6, 7, 8, 9] and time delays [5, 10]. In contrast to prior works such as [5, 6] – which primarily focus on the so-called Smith model – we investigate the robustness of the KLD-RL model. The Smith model is known to be optimal under idealized conditions, such as noise-free payoff observations and infinite populations [11]. In contrast, our study suggests that KLD-RL is better suited for more realistic settings, where noise, modeling errors, and finite agent populations must be taken into account.
@@
Notation: Given a positive integer $n$, let $\mathbb{R}^{n}$ denote the spaces of $n$-dimensional real vectors, $\mathbb{R}_{\geq 0}^{n}$ the space of $n$-dimensional vectors with nonnegative entries, and $\mathbb{R}_{\geq 0}$ the set of nonnegative real numbers. For notational convenience, given positive integers $n$ and $N$, we define the finite set $\mathbb{X}^{N}=\{(X_{1}^{N},\cdots,X_{n}^{N})\in\{0,1/N,\cdots,1\}^{n}\,|\,\sum_{i=1}^{n}X_{i}^{N}=1\}$. Its infinite counterpart is the continuous simplex, defined as $\mathbb{X}=\{(x_{1},\cdots,x_{n})\in[0,1]^{n}\,|\,\sum_{i=1}^{n}x_{i}=1\}$. Note that $\mathbb{X}^{N}$ is a subset of $\mathbb{X}$ for every positive integer $N$. Throughout the paper, we use the Euclidean norm $\|\cdot\|_{2}$ and the maximum norm $\|\cdot\|_{\infty}$. Additionally, given a vector $a$ and a vector-valued function $\mathcal{A}$, the notation $a_{i}$ and $\mathcal{A}_{i}$ refers to the $i$-th component of $a$ and $\mathcal{A}$, respectively.
@@ II-A Evolutionary Dynamics and Task Allocation Games
Consider a finite population of $N$ agents engaged in a multi-agent game [1]. Let the population state at time $t\geq 0$ be represented by the vector $X^{N}(t)=(X^{N}_{1}(t),\cdots,X^{N}_{n}(t))\in\mathbb{X}^{N}$, where each component $X_{i}^{N}(t)$ denotes the fraction of agents selecting the $i$-th strategy among $n$ available options. The agents’ strategy choices determine the payoff vector $p(t)=(p_{1}(t),\cdots,p_{n}(t))\in\mathbb{R}^{n}$, according to an underlying game model, where each $p_{i}(t)$ represents the payoff received by agents selecting strategy $i$. Due to the finite population size, each time an agent switches its strategy, the population state $X^{N}(t)$ undergoes a discrete jump to a new state in the finite set $\mathbb{X}^{N}$. As illustrated in Fig. 1, the payoff vector and population state are jointly determined by a closed-loop system composed of the game dynamics and the decision-making model.
@@
$\rho_{ji}^{\text{\tiny Smith}}(p,X^{N})=\begin{cases}\varrho[p_{i}-p_{j}]_{+}&\text{if }i\neq j\\ 1-\sum_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\varrho[p_{i}-p_{j}]_{+}&\text{otherwise},\\ \end{cases}$ (1)
@@
where $e_{i}$ is the $i$-th standard basis vector, and we use the shorthand notation $\rho_{ij}(t)=\rho_{ij}(p(t),X^{N}(t))$. As shown in [1, Chapter 10], the stochastic process $X^{N}(t)$ can be approximated by the deterministic trajectory $x(t)\in\mathbb{X}$ governed by the following evolutionary dynamics model (EDM):
@@
where $\lambda$ is the rate parameter of the Poisson process, representing the frequency at which each agent revises its strategy.
The accuracy of this approximation improves as the number of agents $N$ increases. In particular, as shown in [12, Lemma 1], when $N$ is sufficiently large and the payoff vector $p(t)$ is a function of $X^{N}(t)$, e.g., $p(t)=F(X^{N}(t))$, the following probabilistic bound holds:
@@
For simplicity, we consider the case where the number of tasks matches the number of available strategies ($m=n$) and assume that the payoff function in (7b) takes the form $G_{i}(q(t),X^{N}(t))=q_{i}(t)$. This assumption simplifies the presentation while preserving the core intuition – agents are more likely to select strategies associated with tasks that have higher job demand. [^1] For a treatment of more general formulations, we refer the interested reader to [2]. We illustrate this model with the following example.
@@
As demonstrated in [3], the model described in (7) can be applied to a multi-robot resource collection scenario. In this setting, $n$ spatially distributed patches contain resources, with $q_{i}(t)$ denoting the volume of remaining resources at patch $i$. The function $\mathcal{F}_{i}$ represents the rate at which agents collect resources from patch $i$ and is defined as:
@@
where $R_{i},\alpha_{i}>0$ and $10<\beta_{i}<$ 1 are model parameters capturing saturation effects and diminishing returns.
@@ II-B Passivity Techniques for Stability Analysis
The concept of passivity [13] has been widely adopted in the game theory literature [14, 15, 16, 17] to study the stability of agent decision-making processes in multi-agent games. Particularly relevant to our work are the notions of $\delta$-passivity and $\delta$-antipassivity, which apply to EDM (5) and the game model (7), respectively [14]. Notably, EDM (5) under the KLD-RL revision protocol (2) satisfies $\delta$-passivity with a positive surplus $\eta^{\ast}>0$, as formally defined below. In contrast, the Smith protocol (1) satisfies $\delta$-passivity without surplus, i.e., $\eta^{\ast}=0$.
@@
$\mathcal{S}(p(t),x(t))-\mathcal{S}(p(t_{0}),x(t_{0}))\leq\int_{\tau=t_{0}}^{t}\big{(}\lambda^{-1}\dot{p}^{\prime}(\tau)\dot{x}(\tau)$
@@
Note that EDM with the KDL-RL protocol satisfies inequality (1) with $\eta^{\ast}=\eta$. In parallel, we define $\delta$-antipassivity for the game model (7) as follows.
@@
$\mathcal{L}(q(t),x(t))\!-\!\mathcal{L}(q(t_{0}),x(t_{0}))\!\leq\!-\!\int_{\tau=t_{0}}^{t}\!\dot{p}^{\prime}(\tau)\dot{x}(\tau)\,\mathrm{d}\tau.$ (10)
As shown in [2], including the case of Example 1, the function $\mathcal{L}$ can be explicitly constructed for a class of task allocation games. Throughout this work, we assume that the game model (7) is $\delta$-antipassive. Furthermore, as discussed in [14], under the assumption of an infinite population of agents, i.e., $N$ is arbitrarily large, the feedback interconnection between a $\delta$-passive model (5) and a $\delta$-antipassive model (7) guarantees the stability of the resulting closed-loop model. In particular, this interconnection guarantees convergence of $(q(t),x(t))$ to an equilibrium of the closed-loop model (see [14, Theorems 4,5] for a detailed discussion).
@@ II-C Problem Formulation
In this work, we consider a setting where: 1)1)1 ) agents maintain an estimate $\hat{p}(t)\in\mathbb{R}^{n}$ of the payoff vector [^2] For simplicity of presentation, we use $\hat{p}(t)$ to denote a common payoff vector estimate shared by all agents. However, as demonstrated in Section III and formally established in [6], our analysis also extends to more general settings in which individual agents maintain distinct payoff vector estimates.; 2)2)2 ) strategy revisions are subject to a constant time delay $d>0$; and 3)3)3 ) the population consists of a finite number of agents, implying that $X^{N}(t)$ and $x(t)$ are not identical, but may be related via the approximation in (6). As a consequence, EDM (5) does not precisely describe the trajectory of $X^{N}(t)$, and the game state $q(t)$ in (7a) evolves based directly on $X^{N}(t)$.
@@
where $\widehat{X}_{i}^{N}(t)$ is the linear interpolation of $X_{i}^{N}(t)$ and $\epsilon_{i}^{N}(t)$ denotes an associated approximation error (as described in Appendix -A). The term $C_{i}^{\eta,\theta}(\hat{p}(t-d))$ represents the delayed strategy selection, incorporating a constant time delay $d$ as formulated in [5]. We can equivalently express (11) as follows:
@@
In our analysis, we focus on the case where the game state $q(t)$ remains bounded and (12) admits a unique equilibrium in the absence of noise. If $\|q(t)\|_{\infty}$ grows unbounded, it indicates that the decision-making model fails to effectively control the task demand. We formalize this as following assumption:
@@
The key research question we address is: How can the parameters $\lambda$ and $\eta$ be selected to minimize $\limsup_{t\to\infty}\|q(t)\|_{\infty}$ – that is, to minimize the maximum remaining job demand across all tasks in the long run? In Section IV, we analyze how these two parameters influence the behavior of the closed-loop model (12). Before delving into the analysis, we first present a simulation example to illustrate the main insights from our theoretical results.
@@ III Illustrative Example
where $\mathbb{N}_{k}$ denotes the set of neighbors of agent $k$ in the underlying communication graph. The communication graph is constructed as a strongly connected Erdős-Rényi random graph with a connection probability of $p=0.2$. One-tenth of the agents are randomly selected as observers; for example, in a simulation with $N=10$ agents, a single agent is designated as an observer with access to $p(t)$. All agents initialize their payoff vector estimates as $\hat{p}^{(k)}(0)=(0,0,0)$.
Simulations are conducted with a population size of $N=10,20,40$. The initial condition is set to $q(0)=(100,200,300)$, and each agent’s initial strategy is selected uniformly at random from the three available strategies. Figure 2 presents the simulation results under various parameter settings. As shown in Fig. 2, the model governed by the KLD-RL protocol (II-A) with $\theta=x^{\ast}$ demonstrates greater robustness to noise compared to the Smith protocol (1) with $\varrho=1/(n-1)M_{q}$, indicating that the KLD-RL protocol is better suited to our problem formulation. Figure 2 illustrates that increasing $\eta$ while keeping $\lambda$ fixed in the KLD-RL protocol reduces $\limsup_{t\to\infty}\|q(t)\|_{\infty}$. However, overly large values of $\eta$ tend to induce increased variance in the trajectory of $\|q(t)\|_{\infty}$. Similarly, Fig. 2 shows that decreasing $\lambda$ in the KLD-RL protocol yields a similar trade-off. Additionally, Fig. 2 demonstrates that the overall performance of the closed-loop – in terms of achieving lower values of $\limsup_{t\to\infty}\|q(t)\|_{\infty}$ and reduced variance – improves as the population size $N$ increases.

### v2

# Robust Decision-Making in Finite-Population Games

@@ Abstract
We study the robustness of an agent decision-making model in finite-population games, with a particular focus on the Kullback-Leibler Divergence Regularized Learning (KLD-RL) model. Specifically, we examine how the model’s parameters influence the impact of various sources of noise and modeling inaccuracies—factors commonly encountered in engineering applications of population games—on agents’ decision-making. Our analysis provides insights into how these parameters can be effectively tuned to mitigate such effects. Theoretical results are supported by numerical examples and simulation studies that validate the analysis and illustrate practical strategies for parameter selection.
@@ I Introduction
The population game and evolutionary dynamics framework provides a powerful foundation for modeling and analyzing repeated strategic interactions among a population of decision-making agents [Sandholm2010-SANPGA-2]. In this work, we consider a class of population games in which agents’ payoffs—driving their strategy selection—are determined by a dynamic payoff mechanism. Specifically, payoff vectors evolve according to a dynamical system model that depends on both the internal state of the game and the population’s state, which reflects agents’ strategy selections. Our primary focus is on a particular subclass of these games, known as task allocation games [10383344, 9561809], where agents interact strategically to accomplish a given set of tasks.
The conventional framework relies on several key assumptions: $1)$ payoff observations are noise-free, $2)$ switching strategies incurs no delay, and $3)$ the population size is infinite. However, these assumptions are often unrealistic in engineering applications of population games. We specifically examine decision-making processes in finite-population games, where agents select strategies based on noisy payoff estimates influenced by their own past actions, and where strategy updates are subject to time delays. A central challenge posed by noise and modeling errors is the increased variability in agents’ decision-making, which can hinder convergence to the optimal strategy selection.
@@
This work builds upon and extends recent studies in multi-agent games [10156113, 10886571, 9993385, 5137415, BRAVO201741, Obando2016], which analyze the impact of noise in payoff observations [10886571, 9993385, 5137415, BRAVO201741] and time delays [10156113, Obando2016]. Unlike our earlier works [10156113, 10886571], which focus primarily on the Smith model, here we examine the robustness of the KLD-RL model. As discussed in Section II, this model exhibits a strong notion of passivity, allowing agents to maintain robustness in the presence of perturbations. Our results therefore suggest that KLD-RL is particularly well-suited for realistic environments, where noise, modeling errors, and finite agent populations play a critical role.
Meanwhile, the conventional population game literature often investigates the perturbed best response model, in which payoffs are subject to random noise [52177154-daeb-3f75-b380-9f9a398bed3c, HOFBAUER200747]. In contrast to our study, these works assume that the noise affecting the payoff vector is independent of both the payoffs themselves and the population state, which reflects the agents’ strategy choices. Our framework, by comparison, explicitly accounts for noise and modeling inaccuracies that depend on these variables.
@@
Notation: Given a positive integer $n$, let $\mathbb{R}^{n}$ denote the spaces of $n$-dimensional real vectors, $\mathbb{R}_{\geq 0}^{n}$ the space of $n$-dimensional vectors with nonnegative entries, and $\mathbb{R}_{\geq 0}$ the set of nonnegative real numbers. For notational convenience, given positive integers $n$ and $N$, we define the finite set $\mathbb{X}^{N}=\{(X_{1}^{N},\cdots,X_{n}^{N})\in\{0,1/N,\cdots,1\}^{n}\,|\,\sum_{i=1}^{n}X_{i}^{N}=1\}$. Its infinite counterpart, obtained as $N\to\infty$, is the continuous simplex $\mathbb{X}=\{(x_{1},\cdots,x_{n})\in[0,1]^{n}\,|\,\sum_{i=1}^{n}x_{i}=1\}$. Note that $\mathbb{X}^{N}$ is a subset of $\mathbb{X}$ for every positive integer $N$. Throughout the paper, we use the Euclidean norm $\|\cdot\|_{2}$ and the maximum norm $\|\cdot\|_{\infty}$. Additionally, given a vector $a$ and a vector-valued function $\mathcal{A}$, the notation $a_{i}$ and $\mathcal{A}_{i}$ refers to the $i$-th component of $a$ and $\mathcal{A}$, respectively.
@@ II-A Evolutionary Dynamics and Task Allocation Games
Consider a finite population of $N$ agents engaged in a multi-agent game [Sandholm2010-SANPGA-2]. Let the population state at time $t\geq 0$ be represented by the vector $X^{N}(t)=(X^{N}_{1}(t),\cdots,X^{N}_{n}(t))\in\mathbb{X}^{N}$, where each component $X_{i}^{N}(t)$ denotes the fraction of agents selecting the $i$-th strategy among $n$ available options. The agents’ strategy choices determine the payoff vector $p(t)=(p_{1}(t),\cdots,p_{n}(t))\in\mathbb{R}^{n}$, according to an underlying game model, where each $p_{i}(t)$ represents the payoff received by agents selecting strategy $i$. Due to the finite population size, each time an agent switches its strategy, the population state $X^{N}(t)$ undergoes a discrete jump to a new state in the finite set $\mathbb{X}^{N}$. As illustrated in Fig. 1, the payoff vector and population state are jointly determined by a closed-loop system composed of the game dynamics and the decision-making model.
@@
[…]
@@
where $e_{i}$ is the $i$-th standard basis vector, and we use the shorthand notation $\rho_{ij}(t)=\rho_{ij}(p(t),X^{N}(t))$. As shown in [Sandholm2010-SANPGA-2, Chapter 10], the stochastic process $X^{N}(t)$ can be approximated by the deterministic trajectory $x(t)\in\mathbb{X}$ governed by the following evolutionary dynamics model (EDM):
@@
where $\lambda$ is the rate parameter of the Poisson process, denoting the frequency at which each agent revises its strategy.
The accuracy of this approximation improves as the number of agents $N$ increases. In particular, as shown in [6fda8e46-15d6-340f-bf79-5c3a826460f2, Lemma 1], when $N$ is sufficiently large and the payoff vector $p(t)$ is a function of $X^{N}(t)$, e.g., $p(t)=F(X^{N}(t))$, the following probabilistic bound holds:
@@
For simplicity, we consider the case where the number of tasks matches the number of available strategies ($m=n$) and assume that the payoff function in (7b) takes the form $G_{i}(q(t),X^{N}(t))=q_{i}(t)$. This assumption simplifies the presentation while preserving the core intuition—agents are more likely to select strategies associated with tasks that have higher job demand. [^1] For more general formulations, we refer the interested reader to [10383344]. We illustrate this model with the following example.
@@
As demonstrated in [9561809], the model described in (7) can be applied to a multi-robot resource collection scenario. In this setting, $n$ spatially distributed patches contain resources, with $q_{i}(t)$ denoting the volume of remaining resources at patch $i$. The function $\mathcal{F}_{i}$ denotes the rate at which agents collect resources from patch $i$ and is given as:
@@
where $R_{i},\alpha_{i}>0$ and $0<\beta_{i}<1$ are model parameters capturing saturation effects and diminishing returns.
@@ II-B Passivity Techniques for Stability Analysis
The concept of passivity [schaft_springer] has been widely adopted in the game theory literature [9029756, 9781277, 9219202, Fox2013Population-Game] to study the stability of agent decision-making processes in multi-agent games. Particularly relevant to our work are the notions of $\delta$-passivity and $\delta$-antipassivity, which apply to EDM (5) and the game model (7), respectively [9029756]. [^2] The symbol $\delta$ in $\delta$-passivity and $\delta$-antipassivity, as introduced in [Fox2013Population-Game], denotes differential passivity in both the EDM and the task allocation game models. Notably, EDM (5) under the KLD-RL revision protocol (2) satisfies $\delta$-passivity with a positive surplus $\eta^{\ast}>0$, as formally defined below. In contrast, the Smith protocol (1) satisfies $\delta$-passivity without surplus, i.e., $\eta^{\ast}=0$.
@@
$\mathcal{S}(p(t),x(t))-\mathcal{S}(p(t_{0}),x(t_{0}))\leq\textstyle\int_{\tau=t_{0}}^{t}\big(\lambda^{-1}\dot{p}^{\prime}(\tau)\dot{x}(\tau)$
@@
Note that EDM with the KLD-RL protocol satisfies inequality (1) with $\eta^{\ast}=\eta$. In parallel, we define $\delta$-antipassivity for the game model (7) as follows.
@@
$\textstyle\mathcal{L}(q(t),x(t))\!-\!\mathcal{L}(q(t_{0}),x(t_{0}))\!\leq\!-\!\int_{\tau=t_{0}}^{t}\!\dot{p}^{\prime}(\tau)\dot{x}(\tau)\,\mathrm{d}\tau.$ (10)
As shown in [10383344], including the case of Example 1, the function $\mathcal{L}$ can be explicitly constructed for a class of task allocation games. Throughout this work, we assume that the game model (7) is $\delta$-antipassive. Furthermore, as discussed in [9029756], under the assumption of an infinite population of agents, i.e., $N$ is arbitrarily large, the feedback interconnection between a $\delta$-passive model (5) and a $\delta$-antipassive model (7) guarantees the stability of the resulting closed-loop model. In particular, this interconnection guarantees convergence of $(q(t),x(t))$ to an equilibrium of the closed-loop model (see [9029756, Theorems 4,5] for a detailed discussion).
@@ II-C Problem Formulation
In this work, we consider a setting where: $1)$ agents maintain an estimate $\hat{p}(t)\in\mathbb{R}^{n}$ of the payoff vector [^3] For simplicity of presentation, we use $\hat{p}(t)$ to denote a common payoff vector estimate shared by all agents. However, as demonstrated in Section III and formally established in [10886571], our analysis also extends to more general settings in which individual agents maintain distinct payoff vector estimates.; $2)$ strategy revisions are subject to a constant time delay $d>0$; and $3)$ the population consists of a finite number of agents, implying that $X^{N}(t)$ and $x(t)$ are not identical, but may be related via the approximation in (6). As a consequence, EDM (5) does not precisely describe the trajectory of $X^{N}(t)$, and the game state $q(t)$ in (7a) evolves based directly on $X^{N}(t)$.
@@
where $\widehat{X}_{i}^{N}(t)$ is the linear interpolation of $X_{i}^{N}(t)$ and $\epsilon_{i}^{N}(t)$ denotes an associated approximation error (as described in Appendix -A). The term $C_{i}^{\eta,\theta}(\hat{p}(t-d))$ denotes the delayed strategy selection, incorporating a constant time delay $d$ as formulated in [10156113]. We can equivalently express (11) as follows:
@@
In our analysis, we focus on the case where the game state $q(t)$ remains bounded and (12) admits a unique equilibrium in the absence of noise. We formalize this as following assumption:
@@
The key research question we address is: How can the parameters $\lambda$ and $\eta$ be selected to minimize $\limsup_{t\to\infty}\|q(t)\|_{\infty}$—that is, to minimize the maximum remaining job demand across all tasks in the long run? In Section IV, we analyze how these two parameters influence the behavior of the closed-loop model (12). Before delving into the analysis, we first present a simulation example to illustrate the main insights from our theoretical results.
Remark 1 (On Assumption 1)
We assume that the equilibrium point $x^{\ast}$ is known. In practice, this information may be unavailable without explicit knowledge of the game model (7) or access to prior data. Unlike the Smith protocol, which serves as our baseline for performance comparison, the KLD-RL protocol (II-A) requires a specified reference distribution $\theta$ in order to realize its full benefits. To assess the potential of the KLD-RL approach, we therefore assume in this paper that $x^{\ast}$ is known a priori.
@@ III Illustrative Example
where $\mathbb{N}_{k}$ denotes the set of neighbors of agent $k$ in the underlying communication graph. The communication graph is constructed as a strongly connected Erdős-Rényi random graph with a connection probability of $0.2$. One-tenth of the agents are randomly selected as observers; for example, in a simulation with $N=10$ agents, a single agent is designated as an observer with access to $p(t)$. All agents initialize their payoff vector estimates as $\hat{p}^{(k)}(0)=(0,0,0)$.
Simulations are conducted with a population size of $N=10,20,40$. The initial condition is set to $q(0)=(100,200,300)$, and each agent’s initial strategy is selected uniformly at random from the three available strategies. Figure 2 presents the simulation results under various parameter settings. As shown in Fig. 2, where the parameter $\lambda$ for the Smith protocol is chosen to minimize $\limsup_{t\to\infty}\|q(t)\|_{\infty}$ over a range of candidate values, the KLD-RL protocol (II-A) with $\theta=x^{\ast}$ demonstrates greater robustness to noise compared to the Smith protocol (1) with $\varrho=1/(n-1)M_{q}$, indicating that the KLD-RL protocol is better suited to our problem formulation. Figure 2 illustrates that increasing $\eta$ while keeping $\lambda$ fixed in the KLD-RL protocol reduces $\limsup_{t\to\infty}\|q(t)\|_{\infty}$. However, overly large values of $\eta$ tend to induce increased variance in the trajectory of $\|q(t)\|_{\infty}$. Similarly, Fig. 2 shows that decreasing $\lambda$ in the KLD-RL protocol yields a similar trade-off. Additionally, Fig. 2 demonstrates that the overall performance of the closed-loop—in terms of achieving lower values of $\limsup_{t\to\infty}\|q(t)\|_{\infty}$ and reduced variance—improves as the population size $N$ increases.